OpenStudy (resoundinggrace):

Medal for you!! Can you help? Solve each proportion 8/x+9=2/x-3

OpenStudy (photon336):

So for this we need to do a number of steps the first is that we multiply by the LCD which is we multiply both by the common denominator which is (x+9)(x-3) \[\frac{ 8 }{ (x+9) } = \frac{ 2 }{ (x-3) }\] \]

OpenStudy (photon336):

\[(x+9)(x-3)\frac{ 8 }{ (x+9) } = \frac{ 2 }{ (x-3) }*(x-3)(x+9)\] now this gives us the following: \[8(x-3) = 2(x+9)\]

OpenStudy (photon336):

I won't go any further than this but you have to distribute the 8 in the ( ) and the same for 2 ( ) then isolate x and solve.

OpenStudy (resoundinggrace):

Awesome thank you so answer is 7 right?

OpenStudy (resoundinggrace):

Can you help with one more?

OpenStudy (resoundinggrace):

Need help on # 25. Have no clue how to figure out.